# Precise Constraints on the Energy Budget of WASP-121 b from its JWST NIRISS/SOSS Phase Curve

### v1

@@ 3.3 Comparison to Bell_EBM
We compare our observations to a simple, semi-analytical energy balance model, Bell_EBM, which accounts for the effects of H2 dissociation (Bell & Cowan, 2018). This process effectively increases the nightside temperature, thereby reducing the day–night temperature contrast. The model assumes there is a single, fully mixed atmospheric layer which absorbs all of the incident radiation and which flows zonally at a constant velocity. This toy model has been previously used to fit phase curve observations (e.g., Dang et al., 2022; Davenport et al., 2025). Comparing to the effective temperature, the EBM should always be fit to the bolometric light curve.
We construct a precomputed grid of energy balance models tailored to the planet. The model includes the option to invert a bolometric light curve into effective temperatures, allowing us to directly compare our observed effective temperatures to the predictions of Bell_EBM.Our grid is generated by fixing the system parameters to the values from Sing et al. (2024) with the exception of the orbital period and inclination which we take from our Order 1 white light curve fit.
@@ 3.4 Reflected and Thermal Components
We present results from the Order 2 detrended white light curve [0.6–0.85 m, effective m] to the slice model described in Section 2.4. Our fits, which account for both thermal and reflected light, indicate that thermal emission dominates the observed flux. In fact, thermal emission alone is sufficient to explain the data. While reflected light can contribute to the flux observed near the dayside, it cannot account for the non-zero flux around transit as the planet reflects virtually no light towards the observer at these phases. Nonetheless, we include results considering both thermal emission and reflected light (Figure 6) to properly marginalize over the full range of possibilities and constrain the amount of reflected light.
@@
We present the best-fit model using the median retrieved parameter values. From the retrieved albedo values we follow Schwartz & Cowan (2015) and infer the geometric albedo as
@@
From this, we measure a geometric albedo of or a 3 upper limit of 0.175. This is consistent with the geometric albedo of found at a similar bandpass in TESS [0.6–0.95 ] (Daylan et al., 2021). Similarly, an eclipse in the z′ band (central wavelength of 0.9 ) suggests a geometric albedo of (Mallonn et al., 2019). Nevertheless, our data provide evidence of reflected light in Order 2.
@@
We compare the median reflected contribution to a Lambertian reflector in Figure 7. The reflected light component shows a distribution that is asymmetric and shifted eastward, suggesting a slightly more reflective eastern dayside of the planet. This is in apparent contradiction with the expectation of inhomogeneous cloud cover which results in greater reflectivity west of the sub-stellar point (Parmentier et al., 2016; Coulombe et al., 2025).
We tested our slice model using 4, 6, and 8 longitudinal slices, allowing albedo values to go negative with a lower prior limit of . One consistent trend emerges across all tests: eastward reflection is always favored. For slices closest to the sub-stellar point, the eastern slice consistently exhibits a higher albedo than the western slice.
@@ 4.1 Observational Systematics and Fit Discrepancies
To test the robustness of our results, we analyzed the data using two independent reduction and fitting pipelines. Our fiducial analysis, performed with exoTEDRF, incorporates physically motivated components such as ellipsoidal variations, a stellar variability model in the Order 2 wavelengths, and a prior punishing negative planetary flux. In contrast, the NAMELESS pipeline omits ellipsoidal variations and stellar variability, and allows negative planetary flux solutions with no constraints. The most significant differences between the two approaches emerge in the inferred nightside temperature and phase offsets, which we explore in detail in the following sub-sections. Importantly, when the exoTEDRF fitting framework is applied to the NAMELESS reduction, the results are consistent with our fiducial analysis, demonstrating that the observed differences arise from the modeling assumptions rather than the data reduction.
From our white light curve fits, if the Gaussian Process model described in Section 2.3.3 corresponds to stellar granulation, then the stellar granulation time scale is minutes for Order 1, and minutes for Order 2. The corresponding stellar granulation amplitudes are ppm (Order 1) and ppm (Order 2), which are consistent with the expected amplitude of 96 ppm (Gilliland et al., 2011). However, the measured granulation time scales are significantly longer than the expected value of 6 minutes. A similar discrepancy was reported by Coulombe et al. (2025).
Increasing the stellar effective temperature from (Delrez et al., 2016) to (Sing et al., 2024) has a slight impact on the calculation of the brightness temperatures, which in turn increases our estimates for the planet’s effective temperature by approximately 30 K. This has a minimal impact on estimations of the planet’s day and nightside temperatures. The more important effect is for the planet’s irradiation temperature. In the analytic model of Cowan & Agol (2011a), increasing the stellar temperature shifts the Bond albedo to higher values for a given planet/star contrast (Equation 20). Therefore, we update the Bond albedo and heat recirculation from previous phase curve observations using the new inferred irradiation temperature.
@@ 4.2 Atmospheric Energy Transport
This discrepancy is driven almost entirely by differences in the inferred nightside effective temperature. In the exoTEDRF fit, the penalty on negative planetary flux allows a broader range of wavelengths to contribute to the nightside temperature estimate, resulting in a narrower and systematically warmer distribution. In contrast, NAMELESS permits negative planetary flux values with no penalties, which are excluded from contributing during the Monte Carlo estimation of the effective temperature. As a result, fewer data points inform the nightside temperature, broadening the distribution and biasing it toward cooler values (see Fig. A6). Although both fits return consistent dayside temperatures, the 120 K offset in nightside temperature leads to significantly different Bond albedo and heat recirculation efficiency estimates. These results underscore how methodological choices can dominate the interpretation of a planet’s atmospheric energy transport, underscoring the critical role of nightside flux modeling in phase‑curve analyses.
@@
Comparing the phase variations in effective temperature to a grid of BELL_EBM energy balance models, we find the best models in the grid prefer a Bond albedo of 0.31, along with low wind speeds of 0.20 km/s and a mixed layer depth near 1 bar. This suggests the EBM favors slightly higher Bond albedos than those inferred solely from dayside and nightside temperatures. However, the Bond albedos remain broadly consistent between the two analyses. The energy balance models also prefer a mixed layer depth pressure of 1 bar: higher regions of the atmosphere respond to the time-variable stellar flux, while deeper layers do not. This is consistent with theoretical expectations for hot Jupiter atmosphere’s photon deposition layer near 1 bar (Parmentier et al., 2013; Tan & Komacek, 2019). This result contrasts with BELL_EBM fits to WASP-121 b’s Spitzer 3.6 and 4.5 m phase curves (Davenport et al., 2025), which inferred mixed layer depths of less than 0.2 bar. The deeper mixed layer found in our analysis is more consistent with expectations for hot Jupiters, though differences in wavelength coverage and observational constraints likely contribute to these variations.
@@
Our BELL_EBM fits favor wind speeds of approximately 200 m/s. This is larger than the sub-50 m/s values found in BELL_EBM fits to WASP-121 b’s Spitzer 3.6 and 4.5 m phase curves (Davenport et al., 2025) but slower than the global zonally averaged wind speeds of 400–500 m/s predicted by general circulation models (Davenport et al., 2025). However, our inferred wind speed is not directly comparable to equatorial jet speeds measured in previous studies (Sing et al., 2024; Seidel et al., 2025). First, BELL_EBM assumes a single global wind speed throughout the atmosphere, whereas high-resolution Doppler spectroscopy has revealed a complex vertical structure in WASP-121 b’s atmospheric circulation. Seidel et al. (2025) measured wind speeds of 4–10 km/s from neutral iron lines tracing lower pressures ( bar). If our inferred mixed layer depth of 1 bar is accurate, then lower wind speeds at these pressures are plausible, with much faster jet streams and day-to-night winds confined to the upper atmosphere. Additionally, JWST NIRSpec observations reveal a wavelength-dependent radial velocity amplitude between NRS1 and NRS2, indicative of average zonal winds of 5.2 km/s. However, wind speeds inferred from BELL_EBM are more analogous to the planet’s bolometric energy transport, likely probing deeper layers than these high-altitude tracers. The inferred slow wind speeds could be a consequence of magnetic drag, in which the planet’s magnetic field impedes atmospheric flow and reduces heat transport efficiency in WASP-121 b (Perna et al., 2010; Menou, 2012).
@@
Hotter planets are generally expected to have shorter radiative timescales resulting in larger day-night temperature contrasts and, therefore, higher phase curve amplitudes (Cowan & Agol, 2011a; Komacek & Showman, 2016). However, observations to date do not consistently show this trend (Dobbs-Dixon & Cowan, 2017; Crossfield, 2015), possibly because phase curves often captured only a fraction of a planet’s total bolometric flux. This apparent discrepancy has motivated further theoretical work exploring other factors that may influence phase curve behavior. For instance, ultra-hot Jupiters typically show smaller phase offsets due to their short radiative timescales (Cowan & Agol, 2011b). Additional processes—such as magnetic drag (Beltz et al., 2022; Beltz & Rauscher, 2024) and nightside clouds (Komacek et al., 2022; Parmentier et al., 2021)—can further shape atmospheric dynamics, enhancing amplitudes and reducing offsets. Conversely, H2 dissociation and recombination may act to redistribute heat more efficiently across the planet, potentially reducing amplitude and increasing the offset (Bell & Cowan, 2018; Tan et al., 2024).
@@
Another factor to consider is the comparison between different observed bandpasses. General circulation models (GCMs), which incorporate more complex physics and atmospheric mixing, generally offer more realistic predictions than simpler semi-analytic models. However, GCMs without atmospheric drag tend to overpredict phase offsets for WASP-121 b (e.g., Parmentier et al., 2018). Recent observations and GCM studies now suggest that WASP-121 b is indeed experiencing atmospheric drag (Wardenier et al., 2024).
@@
Phase offsets of WASP-121 b have been all over the map. Measurements with TESS [0.6–0.95 m] for WASP-121 b show a slight westward offset of , consistent with zero(Daylan et al., 2021). Analysis of two Hubble WFC3 phase curves [1.12–1.64 m] reported a broadband eastward offset of with spectroscopic offsets spanning 0–10∘ (Mikal-Evans et al., 2022). A later re-analysis of the same phase curves indicated a shift in hot spot position between the two epochs and favored models with a 30∘ eastward displacement and a hot spot extent of 50∘ (Changeat et al., 2024). The first analysis of the NIRSpec white light curves reported slight eastward offsets of with NRS1 [2.70–3.72 m] and with NRS2 [3.82–5.15 m] (Mikal-Evans et al., 2023). A more recent full analysis confirms these small eastward shifts in the white light curves, reporting ∘ with NRS1 and ∘ with NRS2, and further extends the results to spectroscopic phase curves, which likewise show slight eastward offsets but never exceeding (Evans-Soma et al., 2025). However, analysis of the Spitzer phase curve in similar bandpasses to NRS1 and NRS2 have reported westward offsets. The analysis of Morello et al. (2023) find offsets of (m) and (m), while a reanalysis of m Spitzer phase curves found an offset of for WASP-121 b (Dang et al., 2025). Furthermore, another independent re-analysis of both Spitzer phase curves of WASP-121 b finds offsets consistent with zero, measuring at m and at m (Davenport et al., 2025).
@@
From our spectroscopic fits we find that the longer wavelengths tend to exhibit small phase offsets (eastward offset of 5.1 1.4∘ in Order 1 white light curve) and a range of phase curve relative amplitudes showing less stark day/night temperature contrasts suggesting potential evidence that we are probing different layers of the atmosphere with varying dynamical timescales (see Fig. 4).
@@ 4.3 Reflected Light and Albedo Constraints
If the geometric albedo of WASP-121 b is comparable to the inferred Bond albedo of 20–30%, we would expect reflected light to dominate the planetary flux at the shortest wavelengths of NIRISS/SOSS. Assuming a gray albedo, this transition from reflected light to thermal emission dominance occurs within the Order 2 wavelengths. Fitting for the thermal emission and reflected light components simultaneously is the only way to properly estimate the contribution of both. Indeed, as shown earlier, fitting the detrended Order 2 white light curve for both a thermal and reflection component reveals that reflection likely has a non-negligible impact. Regardless of the number of longitudinal slices fit or whether we allow albedo to be negative, we consistently find that reflected light peaks on the eastern side of the planet.
@@
The peak in reflection on the eastern side of the planet may be influenced by its strong correlation with thermal emission. However, if this eastward reflection is real, it could be further supported by the detection of large phase offsets at the shortest wavelengths. These offsets appear robust across different modeling approaches (Figure A7), with the NAMELESS fit showing even larger offsets. A notable discrepancy emerges between the two spectral orders, with Order 2 exhibiting a sudden shift to lower phase offsets compared to Order 1. This is likely driven by the inclusion of a stellar variability model, which peaks before secondary eclipse (see Figure 1). When the stellar variability model is excluded, the phase offsets between the two orders become more consistent. Despite these differences, the exoTEDRF fits still show consistently large phase offsets at short wavelengths, a striking and unexpected result. We further note that the trend to higher eastern offsets appears as high as 1.4 m, indicating this is likely astrophysical rather than a systematic difference between Order 1 and Order 2. GCM simulations of this planet, including those that account for clouds, do not find a trend of increasingly eastward offsets at shorter wavelength (Frazier et al., Under Review).
@@
It has been known for a decade that hot Jupiters exhibit greater Bond albedos than geometric albedos, the so-called albedo paradox (Schwartz & Cowan, 2015; Crossfield, 2015). The Bond albedo is constrained through thermal emission, while the geometric albedo is typically determined from reflected light at short optical wavelengths. In this work, we constrain both albedos within a single observation by probing the planet’s reflection and emission and across optical and near-infrared wavelengths. Due to the unprecedented captured flux of our measurement, we provide the most precise estimate of Bond albedo for an exoplanet to date. Our results indicate a relatively high Bond albedo of 20-–30% but a low geometric albedo 10%, a factor of three discrepancy. This finding further sharpens the albedo paradox, underscoring the need for future studies to reconcile the reflected light observations of WASP-121 b and other hot Jupiters to better understand the underlying atmospheric processes.

### v2

@@ 3.3 Comparison to Bell_EBM
We compare our observations to a simple, semi-analytical energy balance model, Bell_EBM, which accounts for the effects of H2 dissociation (Bell2018). This process effectively increases the nightside temperature, thereby reducing the day–night temperature contrast. The model assumes there is a single, fully mixed atmospheric layer that absorbs all of the incident radiation and that flows zonally at a constant velocity. This toy model has been previously used to fit phase curve observations (e.g., Dang2022; Davenport2025). Comparing to the effective temperature, the EBM should always be fit to the bolometric light-curve.
We construct a precomputed grid of energy balance models tailored to the planet. The model includes the option to invert a bolometric light-curve into effective temperatures, allowing us to directly compare our observed effective temperatures to the predictions of Bell_EBM. Our grid is generated by fixing the system parameters to the values from Sing2024-NIRSpec with the exception of the orbital period and inclination that we take from our Order 1 white light-curve fit.
@@ 3.4 Reflected and Thermal Components
We present results from the Order 2 detrended white light-curve [0.6–0.85 m, effective m] to the slice model described in Section 2.4. Our fits, which account for both thermal and reflected light, indicate that thermal emission dominates the observed flux. In fact, thermal emission alone is sufficient to explain the data. While reflected light can contribute to the flux observed near the dayside, it cannot account for the nonzero flux around transit as the planet reflects virtually no light toward the observer at these phases. Nonetheless, we include results considering both thermal emission and reflected light (Figure 5) to properly marginalize over the full range of possibilities and constrain the amount of reflected light.
@@
We present the best-fit model using the median retrieved parameter values. From the retrieved albedo values we follow Schwartz2015 and infer the geometric albedo as
@@
From this, we measure a geometric albedo of or a 3 upper limit of 0.175. This is consistent with the geometric albedo of found at a similar bandpass in TESS [0.6–0.95 ] (Daylan2021). Similarly, an eclipse in the z′ band (central wavelength of 0.9 ) suggests a geometric albedo of (Mallonn2019). Nevertheless, our data provide evidence of reflected light in Order 2.
@@
We compare the median reflected contribution to a Lambertian reflector in Figure 6. The reflected light component shows a distribution that is asymmetric and shifted eastward, suggesting a slightly more reflective eastern dayside of the planet. This is in apparent contradiction with the expectation of inhomogeneous cloud cover, which results in greater reflectivity west of the substellar point (Parmentier2016; Coulombe2025-LTT-NIRISS-PC).
We tested our slice model using 4, 6, and 8 longitudinal slices, allowing albedo values to go negative with a lower prior limit of . One consistent trend emerges across all tests: eastward reflection is always favored. For slices closest to the substellar point, the eastern slice consistently exhibits a higher albedo than the western slice.
@@ 4.1 Observational Systematics and Fit Discrepancies
To test the robustness of our results, we analyzed the data using two independent reduction and fitting pipelines. Our fiducial analysis, performed with exoTEDRF, incorporates physically motivated components such as ellipsoidal variations, a stellar variability model in the Order 2 wavelengths, and a prior punishing negative planetary flux. In contrast, the NAMELESS pipeline omits ellipsoidal variations and stellar variability, and allows negative planetary flux solutions with no constraints. The most significant differences between the two approaches emerge in the inferred nightside temperature and phase offsets, which we explore in detail in the following subsections. Importantly, when the exoTEDRF fitting framework is applied to the NAMELESS reduction, the results are consistent with our fiducial analysis, demonstrating that the observed differences arise from the modeling assumptions rather than the data reduction.
From our white light-curve fits, if the Gaussian process model described in Section 2.3.3 corresponds to stellar granulation, then the stellar granulation time scale is minutes for Order 1 and minutes for Order 2. The corresponding stellar granulation amplitudes are ppm (Order 1) and ppm (Order 2), which are consistent with the expected amplitude of 96 ppm (Gilliland2011). However, the measured granulation time scales are significantly longer than the expected value of 6 minutes. A similar discrepancy was reported by Coulombe2025-LTT-NIRISS-PC.
Increasing the stellar effective temperature from (Delrez2016-discovery) to (Sing2024-NIRSpec) has a slight impact on the calculation of the brightness temperatures, which in turn increases our estimates for the planet’s effective temperature by approximately 30 K. This has a minimal impact on estimations of the planet’s day and nightside temperatures. The more important effect is for the planet’s irradiation temperature. In the analytic model of Cowan2011-albedo, increasing the stellar temperature shifts the Bond albedo to higher values for a given planet/star contrast (Equation 20). Therefore, we update the Bond albedo and heat recirculation from previous phase curve observations using the new inferred irradiation temperature.
@@ 4.2 Atmospheric Energy Transport
This discrepancy is driven almost entirely by differences in the inferred nightside effective temperature. In the exoTEDRF fit, the penalty on negative planetary flux allows a broader range of wavelengths to contribute to the nightside temperature estimate, resulting in a narrower and systematically warmer distribution. In contrast, NAMELESS permits negative planetary flux values with no penalties, which are excluded from contributing during the Monte Carlo estimation of the effective temperature. As a result, fewer data points inform the nightside temperature, broadening the distribution and biasing it toward cooler values (see Fig. C6). Although both fits return consistent dayside temperatures, the 120 K offset in nightside temperature leads to significantly different Bond albedo and heat recirculation efficiency estimates. These results underscore how methodological choices can dominate the interpretation of a planet’s atmospheric energy transport, underscoring the critical role of nightside flux modeling in phase‑curve analyses.
@@
Comparing the phase variations in effective temperature to a grid of BELL_EBM energy balance models, we find the best models in the grid prefer a Bond albedo of 0.31, along with low wind speeds of 0.20 km/s and a mixed layer depth near 1 bar (see Figure 7. This suggests the EBM favors slightly higher Bond albedos than those inferred solely from dayside and nightside temperatures. However, the Bond albedos remain broadly consistent between the two analyses. The energy balance models also prefer a mixed layer depth pressure of 1 bar: higher regions of the atmosphere respond to the time-variable stellar flux, while deeper layers do not. This is consistent with theoretical expectations for hot Jupiter atmosphere’s photon deposition layer near 1 bar (Parmentier2013; Tan2019). This result contrasts with BELL_EBM fits to WASP-121 b’s Spitzer 3.6 and 4.5 m phase curves (Davenport2025), which inferred mixed layer depths of less than 0.2 bar. The deeper mixed layer found in our analysis is more consistent with expectations for hot Jupiters, though differences in wavelength coverage and observational constraints likely contribute to these variations.
@@
Our BELL_EBM fits favor wind speeds of approximately 200 m/s. This is larger than the sub-50 m/s values found in BELL_EBM fits to WASP-121 b’s Spitzer 3.6 and 4.5 m phase curves (Davenport2025) but slower than the global zonally averaged wind speeds of 400–500 m/s predicted by general circulation models (Davenport2025). However, our inferred wind speed is not directly comparable to equatorial jet speeds measured in previous studies (Sing2024-NIRSpec; Seidel2025). First, BELL_EBM assumes a single global wind speed throughout the atmosphere, whereas high-resolution Doppler spectroscopy has revealed a complex vertical structure in WASP-121 b’s atmospheric circulation. Seidel2025 measured wind speeds of 4–10 km/s from neutral iron lines tracing lower pressures ( bar). If our inferred mixed layer depth of 1 bar is accurate, then lower wind speeds at these pressures are plausible, with much faster jet streams and day-to-night winds confined to the upper atmosphere. Additionally, JWST NIRSpec observations reveal a wavelength-dependent radial velocity amplitude between NRS1 and NRS2, indicative of average zonal winds of 5.2 km/s. However, wind speeds inferred from BELL_EBM are more analogous to the planet’s bolometric energy transport, likely probing deeper layers than these high-altitude tracers. The inferred slow wind speeds could be a consequence of magnetic drag, in which the planet’s magnetic field impedes atmospheric flow and reduces heat transport efficiency in WASP-121 b (Perna2010; Menou2012).
@@
Hotter planets are generally expected to have shorter radiative timescales resulting in larger day-night temperature contrasts and, therefore, higher phase curve amplitudes (Cowan2011-albedo; Komacek&Showman2016). However, observations to date do not consistently show this trend (Dobbs-Dixon&Cowan2017; Crossfield2015), possibly because phase curves often captured only a fraction of a planet’s total bolometric flux. This apparent discrepancy has motivated further theoretical work exploring other factors that may influence phase curve behavior. For instance, ultra-hot Jupiters typically show smaller phase offsets due to their short radiative timescales (Cowan&Agol-timescales). Additional processes—such as magnetic drag (Beltz2022; Beltz2024) and nightside clouds (Komacek-ultrahotclouds; Parmentier2021)—can further shape atmospheric dynamics, enhancing amplitudes and reducing offsets. Conversely, H2 dissociation and recombination may act to redistribute heat more efficiently across the planet, potentially reducing amplitude and increasing the offset (Bell2018; Tan2024).
@@
Another factor to consider is the comparison between different observed bandpasses. General circulation models (GCMs), which incorporate more complex physics and atmospheric mixing, generally offer more realistic predictions than simpler semi-analytic models. However, GCMs without atmospheric drag tend to overpredict phase offsets for WASP-121 b (e.g., Parmentier2018). Recent observations and GCM studies now suggest that WASP-121 b is indeed experiencing atmospheric drag (Wardenier2024).
@@
Phase offsets of WASP-121 b have been all over the map. Measurements with TESS [0.6–0.95 m] for WASP-121 b show a slight westward offset of , consistent with zero(Daylan2021). Analysis of two Hubble WFC3 phase curves [1.12–1.64 m] reported a broadband eastward offset of with spectroscopic offsets spanning 0–10∘ (Mikal-Evans2022). A later re-analysis of the same phase curves using a three-region model (and therefore not directly comparable to phase offsets) indicated a shift in hot spot position between the two epochs and favored models with a 30∘ eastward displacement and a hot spot extent of 50∘ (Changeat2024). The first analysis of the NIRSpec white light-curves reported slight eastward offsets of with NRS1 [2.70–3.72 m] and with NRS2 [3.82–5.15 m] (Mikal-Evans2023). A more recent full analysis confirms these small eastward shifts in the white light-curves, reporting ∘ with NRS1 and ∘ with NRS2, and further extends the results to spectroscopic phase curves, which likewise show slight eastward offsets but never exceeding (Evans-Soma2025-NIRSpec-PhaseCurve). However, analysis of the Spitzer phase curve in similar bandpasses to NRS1 and NRS2 have reported westward offsets. The analysis of Morello2023 find offsets of (m) and (m), while a reanalysis of m Spitzer phase curves found an offset of for WASP-121 b (Dang2025). Furthermore, another independent re-analysis of both Spitzer phase curves of WASP-121 b finds offsets consistent with zero, measuring at m and at m (Davenport2025).
@@
From our spectroscopic fits we find that the longer wavelengths tend to exhibit small phase offsets (eastward offset of 5.1 1.4∘ in Order 1 white light-curve) and a range of phase curve relative amplitudes showing less stark day/night temperature contrasts suggesting potential evidence that we are probing different layers of the atmosphere with varying dynamical timescales (see Fig. 4).
@@ 4.3 Reflected Light and Albedo Constraints
If the geometric albedo of WASP-121 b is comparable to the inferred Bond albedo of 20–30%, we would expect reflected light to dominate the planetary flux at the shortest wavelengths of NIRISS/SOSS. Assuming a gray albedo, this transition from reflected light to thermal emission dominance occurs within the Order 2 wavelengths. Fitting for the thermal emission and reflected light components simultaneously is the only way to properly estimate the contribution of both. Indeed, as shown earlier, fitting the detrended Order 2 white light-curve for both a thermal and reflection component reveals that reflection likely has a non-negligible impact. Regardless of the number of longitudinal slices fit or whether we allow albedo to be negative, we consistently find that reflected light peaks on the eastern side of the planet.
@@
The peak in reflection on the eastern side of the planet may be influenced by its strong correlation with thermal emission. However, if this eastward reflection is real, it could be further supported by the detection of large phase offsets at the shortest wavelengths. These offsets appear robust across different modeling approaches (Figure C7), with the NAMELESS fit showing even larger offsets. A notable discrepancy emerges between the two spectral orders, with Order 2 exhibiting a sudden shift to lower phase offsets compared to Order 1. This is likely driven by the inclusion of a stellar variability model, which peaks before secondary eclipse (see Figure 1). When the stellar variability model is excluded, the phase offsets between the two orders become more consistent. Despite these differences, the exoTEDRF fits still show consistently large phase offsets at short wavelengths, a striking and unexpected result. We further note that the trend to higher eastern offsets appears as high as 1.4 m, indicating this is likely astrophysical rather than a systematic difference between Order 1 and Order 2. GCM simulations of this planet, including those that account for clouds, do not find a trend of increasingly eastward offsets at shorter wavelength (Frazier et al., Under Review).
@@
It has been known for a decade that hot Jupiters exhibit greater Bond albedos than geometric albedos, the so-called albedo paradox (Schwartz2015; Crossfield2015). The Bond albedo is constrained through thermal emission, while the geometric albedo is typically determined from reflected light at short optical wavelengths. In this work, we constrain both albedos within a single observation by probing the planet’s reflection and emission and across optical and near-infrared wavelengths. Due to the unprecedented captured flux of our measurement, we provide the most precise estimate of Bond albedo for an exoplanet to date. Our results indicate a relatively high Bond albedo of 20-–30% but a low geometric albedo 10%, a factor of 3 discrepancy. This finding further sharpens the albedo paradox, underscoring the need for future studies to reconcile the reflected light observations of WASP-121 b and other hot Jupiters to better understand the underlying atmospheric processes.